# Towards Fair and Efficient Public Transportation: A Bus Stop Model

### v1

@@ 4.1 Theoretical Possibilities and Limitations
Given an instance $\mathcal{I}=\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ of 00-BSP, Algorithm 1 starts by sorting $V=\{v_{1},\dots,v_{m}\}$ so that $v_{1}\leq v_{2}\leq\dots\leq v_{m}$. Then, for each $j\in[m]$, the algorithm computes $x_{j}$ as the number of agent terminals at or to the left of $v_{j}$. Next, for $k\in[b]$, it computes $s_{k}$ as the leftmost element $v_{j}\in V$ with $x_{j}\geq k\left\lfloor\frac{2n}{b}\right\rfloor$; it then returns $S=\{s_{k}:k\in[b]\}$. Since $|S|\leq b$, $S$ is a feasible solution.
@@
Let $\mathcal{I}=\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ be an instance of 00-BSP, and let $S$ be a solution for $\mathcal{I}$. Consider a coalition $M\subseteq N$ that prefers $T\subseteq V$ to $S$, and an agent $i\in M$. Suppose that, when $T$ is built, $i$ walks from their left terminal $\ell_{i}$ to $\ell^{\prime}\in T$, then takes the bus from $\ell^{\prime}$ to $r^{\prime}\in T$, and then walks from $r^{\prime}$ to their right terminal $r_{i}$. If there exists an $\ell\in S$ with $\ell_{i}\leq\ell\leq\ell^{\prime}$ or $\ell^{\prime}\leq\ell\leq\ell_{i}$, then, for every stop $r\in S$ it holds that $d(\ell,\ell^{\prime})<d(r,r^{\prime}).$ Similarly, if there exists a stop $r\in S$ with $r_{i}\leq r\leq r^{\prime}$ or $r^{\prime}\leq r\leq r_{i}$, then, for every stop $\ell\in S$ it holds that $d(\ell,\ell^{\prime})>d(r,r^{\prime}).$
@@
Assume that $M$, $S$, $T$, $i$, $\ell^{\prime}$ and $r^{\prime}$ are as in the statement of the lemma and that there exists an $\ell\in S$ between $\ell_{i}$ and $\ell^{\prime}$. Define $r^{*}:=\operatorname*{arg\,min}_{x\in S}d(r_{i},x)$. By assumption, we have $c_{i}(T)<c_{i}(S)$. Hence, for every $r\in S$, we conclude that
@@
where we first use the minimality of $r^{*}$ and then the triangle inequality. The proof for the second statement of the lemma is analogous. ∎
Algorithm 1 JR solutions for 00-BSP.
Input: Instance $\mathcal{I}=\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ of 00-BSP
Output: A solution $S$ that provides JR
@@
Algorithm 1 runs in polynomial time, and for 00-BSP it computes feasible solutions that provide JR.
@@
This instance is illustrated in Figure 2. Note that agents $1$ and $2$ have symmetric roles with respect to the bus stop at 00. Moreover, the budget is equal to the number of agents.
@@
Now, consider their costs under $S$. Agent $1$ has to start by walking for one step. Then, between $-2-\lambda$ and 00 it is fastest to take the bus. Finally, there are two options for travelling from 00 to $\lambda$, the destination of agent $1$. Taking the bus and walking back has a cost of $\alpha(\lambda+2)+2=2(1+\alpha)+\alpha\lambda$, whereas walking has a cost of $\lambda$. Note that $\lambda>2(1+\alpha)+\alpha\lambda$ if and only if $\lambda>2\frac{1+\alpha}{1-\alpha}$, which is exactly how we chose $\lambda$. Hence, taking the bus and then walking back is faster. Thus, agent $1$ will walk from $-3-\lambda$ to $-2-\lambda$, take the bus to $\lambda+2$ and walk back to $\lambda$, which has a cost of $c_{1}(S)=3+\alpha(2\lambda+4)>c_{1}(T)$. A similar computation shows that $c_{2}(S)=3+\alpha(2\lambda+4)>c_{2}(T)$. Together, this shows that $S$ fails to provide JR. ∎
In contrast, for $\alpha=0$ we can show stronger fairness guarantees for Algorithm 1, namely that it computes solutions in the $2$-core. However, the approximation guarantee of $2$ is tight.
@@
For 00-BSP, Algorithm 1 computes solutions in the $2$-core. However, for each $0<\epsilon\leq 1$, it may output solutions that are not in the $(2-\epsilon)$-core.
@@
We start by proving that the output of Algorithm 1 is in the $2$-core. Consider an instance $\mathcal{I}=\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ and a solution $S$ computed by Algorithm 1 on $\mathcal{I}$. Assume for contradiction that there is a set of agents $M\subseteq N$ and a set of stops $T\subseteq V$ with $|T|\leq\frac{b}{2n}\cdot|M|$ such that $c_{i}(T)<c_{i}(S)$ for all $i\in M$.
For each $t\in T$, let $C_{t}:=\{i\in N\colon d(\ell_{i},t)<d(\ell_{i},S)\text{ or }d(r_{i},t)<d(r_{i},S)\}$, i.e., $C_{t}$ is the set of agents for whom $t$ is closer to one of their terminals than any stop in $S$. If $t\in S$, we have $C_{t}=\varnothing$. On the other hand, if $t$ lies between two adjacent stops $s,s^{\prime}\in S$, each agent in $C_{t}$ has at least one terminal that lies strictly between $s$ and $s^{\prime}$, so by the design of Algorithm 1 we have $|C_{t}|\leq\frac{2n}{b}-1$. By a similar argument, if $t$ lies to the left of the leftmost stop in $S$ or to the right of the rightmost stop in $S$, we also have $|C_{t}|\leq\frac{2n}{b}-1$. Hence,
$\left|\bigcup_{t\in T}C_{t}\right|\leq\sum_{t\in T}|C_{t}|\leq|T|\left(\frac{2n}{b}-1\right)<|M|\text{,}$
so $M\setminus\left(\bigcup_{t\in T}C_{t}\right)\neq\varnothing$. Consider an agent $i^{*}\in M\setminus\left(\bigcup_{t\in T}C_{t}\right)$. By definition of the sets $C_{t}$, there exist stops $\ell^{\prime},r^{\prime}\in S$ with $d(\ell_{i^{*}},\ell^{\prime})\leq d(\ell_{i^{*}},T)$ and $d(r_{i^{*}},r^{\prime})\leq d(r_{i^{*}},T)$. Now, if the cost of $i^{*}$ with respect to $T$ is minimized when walking, it clearly holds that $c_{i^{*}}(S)\leq c_{i^{*}}(T)$. Otherwise, it holds that
$c_{i^{*}}(S)\leq d(\ell_{i^{*}},\ell^{\prime})+d(r_{i^{*}},r^{\prime})\leq d(\ell_{i^{*}},T)+d(r_{i^{*}},T)=c_{i^{*}}(T)\text{.}$
For the last transition, we use that $i^{*}$ is taking the bus and $\alpha=0$. We have derived that $c_{i^{*}}(S)\leq c_{i^{*}}(T)$, a contradiction with $i^{*}\in M$.
To prove that the bound is tight, let $0<\epsilon\leq 1$. We define an instance that is parameterized by an integer $k$. First, observe that $\frac{k+1}{2k}(2-\epsilon)$ converges to $1-\frac{\epsilon}{2}$ as $k$ tends to infinity. Hence, we can choose $k$ large enough so that $\frac{k+1}{2k}(2-\epsilon)\leq 1-\frac{\epsilon}{4}$. Moreover, fix an integer $s>\frac{4}{\epsilon}$.
We define an instance with $b=2k$ and $k$ groups of agents $N^{1},\dots,N^{k}$ of size $s$ each, so $n=ks$. We want the agents’ left and right terminals to be separated by a sufficiently long part of the path with no agent terminals. To this end, for each $j\in[k]$, we let $\ell^{j}=j$, $r^{j}=2k+(2j-1)k$, and set
  • $\ell_{i}=\ell^{j}$ for each $i\in N^{j}$;
  • $r_{i}=r^{j}$ for $s-1$ agents in $N^{j}$ and $r_{i}=r^{j}+k$ for the remaining agent in $N^{j}$.
@@
For $V=\cup_{i\in N}\{\ell_{i},r_{i}\}$, the algorithm outputs the bus stops in $S=[k]\cup\{2k+2jk\colon j\in[k]\}$.
@@
$|M|$ $=n-k=k(s-1)>k\left(s-\frac{\epsilon}{4}s\right)=\left(1-\frac{\epsilon}{4}\right)ks$
$\geq\frac{k+1}{2k}\left(2-\epsilon\right)ks=\left(k+1\right)\left(2-\epsilon\right)\frac{n}{b}\text{.}$
Here, the first inequality holds because of our bound on $s$, and the second inequality holds because of our choice of $k$. Consider the set of bus stops
$T=\{1\}\cup\{2k+(2j-1)k\colon j\in[k]\}.$
Then, $|T|=k+1$. Moreover, for every agent $i\in M$ it holds that $c_{i}(T)\leq k-1<k=c_{i}(S)$. Hence, $M$ is a blocking coalition for the $(2-\epsilon)$-core. ∎
While Theorem 4.4 only applies to the case $\alpha=0$, the existence of solutions in the approximate core is guaranteed for any $\alpha$. For this, we leverage a result by Jiang et al. (2020).
Theorem 4.5.
Let $\alpha\in(0,1)$. Then for $\alpha$-BSP there exists a solution in the $16$-core.
Proof.
Let $\alpha\in(0,1)$. For every agent $i\in N$ and for every pair of subsets of stops $S,T\subseteq V$ with $S\subseteq T$ it holds that $c_{i}(S)\geq c_{i}(T)$. Hence, $\alpha$-BSP can be viewed as a special case of committee selection with monotonic preferences (see Appendix D). Theorem 1 by Jiang et al. (2020) implies that there exists a solution in the $32$-core. Since we have a uniform cost for bus stops, this improves to an approximation factor of $16$. ∎
It remains an open problem how to construct solutions in the core; in fact, we do not even know if the core is always non-empty. This seems to be a challenging question. For instance, by Proposition 2.5, the cost-minimal solution does not even provide JR. Similarly, the solution that selects the most popular terminal points may not provide JR. An example is provided in Appendix A.
@@ 4.2 Computation of Outcomes in the Core on Synthetic Data
We now present evidence from simulations that indicates that Algorithm 1 can frequently find outcomes in the core. This is despite the fact that, according to Proposition 4.3 and Theorem 4.4, in the worst case this algorithm may fail to find JR outcomes for $\alpha\in(0,1)$ and outcomes in the $\beta$-core for $\beta<2$. All computations were performed using an Apple M2 CPU with 24 GB of RAM. [^1] All codes are available at https://github.com/latifian/Bus_Stop_Model.
@@ 4.2.1 Experimental Setup
Figure 5: A heat map showing the frequency of core violations (as a percentage, i.e., 0.05 means 0.05%) of the solutions computed by Algorithm 1 for pairs of $m$ and $b$ in instances of 00-BSP. Each cell is averaged over all values of $n$.
@@
We measure the frequency of instances on which Algorithm 1 computes solutions that provide JR/are in the core. For comparison, we benchmark our algorithm against the naive algorithm, which selects the stops as if the agents were distributed uniformly on the line. More precisely, if $V$ is the set of possible stops, the naive algorithm ignores the actual agents’ types and assumes instead that for each point $p\in\{1,\ldots,100\}$, $n/50$ agents have one of their endpoints at $p$ and report a point in $\arg\min d(p,V)$ as the respective terminal; it then runs Algorithm 1 under this assumption on the agents’ types. In other words, Algorithm 1 can be seen as a weighted version of the naive algorithm, which takes actual user demands into consideration. In Section C.3, we consider a second benchmark algorithm based on the idea of selecting the most demanded bus stops.
@@ 4.2.2 Experimental Results
The primary goal of our experiments is to measure the performance of Algorithm 1 in terms of achieving outcomes in the core or providing JR beyond the guarantee of Theorem 4.2. Figure 5 shows an aggregated view of our results. We see that, for a fixed $\alpha$, Algorithm 1 computes a solution in the core in more than $99.9\%$ of the cases. Even for the parameter combinations with the highest frequency of fairness violations, the failure rate of Algorithm 1 with respect to computing solutions in the core does not exceed 3%. Figure 5 gives a glimpse at a more nuanced distribution of the failures of computing solutions in the core. We defer a more detailed analysis to Section C.2. We remark that the upper left triangle of the figure cannot contain any core violations because then the budget exceeds the number of stops. One trend that be observed from Figure 5 and that is confirmed in our analysis for any fixed number of agents is that the frequency of core violations is highest for a comparatively high number of potential bus stops and a smaller budget.
Another interesting observation is that, for a significant fraction of the instances, if the solution computed by Algorithm 1 is not in the core, it fails JR as well; indeed, this fraction tends to increase as the cost parameter $\alpha$ increases. Moreover, Algorithm 1 performs much better than the naive algorithm. On average it performs up to $230.2$ better for smaller values of $\alpha$ and still $83.7$ times better for $\alpha=0.9$. Detailed numbers comparing their performance are provided in Table 1 in Section C.2.
Our interpretation of this finding is that, while placing bus stops uniformly is a simple and appealing approach, taking into account the actual user demands results in much fairer solutions.

### v2

@@ 4.1 Theoretical Possibilities and Limitations
Given an instance $\mathcal{I}=\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ of 0-BSP, Algorithm 1 starts by sorting $V=\{v_{1},\dots,v_{m}\}$ so that $v_{1}\leq v_{2}\leq\dots\leq v_{m}$. Then, for each $j\in[m]$, the algorithm computes $x_{j}$ as the number of agent terminals at or to the left of $v_{j}$. Next, for $k\in[b]$, it computes $s_{k}$ as the leftmost element $v_{j}\in V$ with $x_{j}\geq k\left\lfloor\frac{2n}{b}\right\rfloor$; it then returns $S=\{s_{k}:k\in[b]\}$. Since $|S|\leq b$, $S$ is a feasible solution.
@@
Let $\mathcal{I}=\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ be an instance of 0-BSP, and let $S$ be a solution for $\mathcal{I}$. Consider a coalition $M\subseteq N$ that prefers $T\subseteq V$ to $S$, and an agent $i\in M$. Suppose that, when $T$ is built, $i$ walks from their left terminal $\ell_{i}$ to $\ell^{\prime}\in T$, then takes the bus from $\ell^{\prime}$ to $r^{\prime}\in T$, and then walks from $r^{\prime}$ to their right terminal $r_{i}$. If there exists an $\ell\in S$ with $\ell_{i}\leq\ell\leq\ell^{\prime}$ or $\ell^{\prime}\leq\ell\leq\ell_{i}$, then for every stop $r\in S$ it holds that $d(\ell,\ell^{\prime})<d(r,r^{\prime}).$ Similarly, if there exists a stop $r\in S$ with $r_{i}\leq r\leq r^{\prime}$ or $r^{\prime}\leq r\leq r_{i}$, then for every stop $\ell\in S$ it holds that $d(\ell,\ell^{\prime})>d(r,r^{\prime}).$
@@
Assume that $M$, $S$, $T$, $i$, $\ell^{\prime}$ and $r^{\prime}$ are as in the statement of the lemma and that there exists an $\ell\in S$ between $\ell_{i}$ and $\ell^{\prime}$. Fix an $r^{*}$ in $\operatorname*{arg\,min}_{x\in S}d(r_{i},x)$. By assumption, we have $c_{i}(T)<c_{i}(S)$. Hence, for every $r\in S$, we conclude that
@@
where the first transition is by the choice of $r^{*}$ and the second transition uses the triangle inequality. The proof for the second statement of the lemma is analogous. ∎
Algorithm 1 Fair solutions for $\alpha$-BSP.
Input: Instance $\mathcal{I}=\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ of $\alpha$-BSP
Output: A solution $S$
@@
Algorithm 1 runs in polynomial time, and for 0-BSP it computes feasible solutions that provide JR.
@@
Example 4.3.
[…]
The example shows that, while Algorithm 1 always returns a feasible solution that provides JR, the obtained solution may fail to exhaust the budget. In such cases, having achieved JR via Algorithm 1, we can distribute the remaining budget to accomplish other goals. For instance, as per Theorem 3.4, we can extend the solution to lower the total cost as much as possible. Note that, after we add stops, the solution continues to provide JR.
@@
This instance is illustrated in Figure 2. Note that agents $1$ and $2$ have symmetric roles with respect to the bus stop at 0. Moreover, the budget is equal to the number of agents.
@@
Now, consider their costs under $S$. Agent $1$ has to start by walking for one step. Then, between $2-\lambda$ and 0 it is fastest to take the bus. Finally, there are two options for traveling from 0 to $\lambda$, the destination of agent $1$. Taking the bus and walking back has a cost of $\alpha(\lambda+2)+2=2(1+\alpha)+\alpha\lambda$, whereas walking has a cost of $\lambda$. Note that $\lambda>2(1+\alpha)+\alpha\lambda$ if and only if $\lambda>2\frac{1+\alpha}{1-\alpha}$, which is exactly how we chose $\lambda$. Hence, taking the bus and then walking back is faster. Thus, agent $1$ will walk from $3-\lambda$ to $2-\lambda$, take the bus to $\lambda+2$ and walk back to $\lambda$, which has a cost of $c_{1}(S)=3+\alpha(2\lambda+4)>c_{1}(T)$. A similar computation shows that $c_{2}(S)=3+\alpha(2\lambda+4)>c_{2}(T)$. Together, this shows that $S$ fails to provide JR. ∎
We can show further fairness guarantees for Algorithm 1, namely that it computes solutions in the $2$-core. However, the approximation guarantee of $2$ is tight.
@@
Let $\alpha\in[0,1)$. Then, for $\alpha$-BSP, Algorithm 1 computes solutions in the $2$-core. However, for each $0<\epsilon\leq 1$, it may output solutions that are not in the $(2-\epsilon)$-core.
@@
Let $\alpha\in[0,1)$. We start by proving that the output of Algorithm 1 is in the $2$-core. Consider an instance $\mathcal{I}=\langle N,V,b,(\theta_{i})_{i\in N}\rangle$ and a solution $S$ computed by Algorithm 1 on $\mathcal{I}$. Assume for contradiction that there is a set of agents $M\subseteq N$ and a set of stops $T\subseteq V$ with $|T|\leq\frac{b}{2n}\cdot|M|$ such that $c_{i}(T)<c_{i}(S)$ for all $i\in M$.
For each $t\in T$, define $\ell^{t}:=\max(\{s\in S\colon s\leq t\}\cup\{-\infty\})$ and $r^{t}:=\min(\{s\in S\colon s\geq t\}\cup\{\infty\})$. These are the closest bus stops in $S$ to the left and to the right of $t$; if there is no such stop, we set this variable to -$\infty$ or $\infty$, respectively. Let $C_{t}:=\{i\in N\colon\ell^{t}<\ell_{i}<r^{t}\text{ or }\ell^{t}<r_{i}<r^{t}\}$, i.e., $C_{t}$ is the set of agents that have at least one terminal strictly between $\ell^{t}$ and $r^{t}$. By design of Algorithm 1, we have
$|C_{t}|\leq\frac{2n}{b}-1\text{.}$ (4)
We claim that for every $i\in M$ there exists $t\in T$ with $i\in C_{t}$. Indeed, since $c_{i}(T)<c_{i}(S)$, we know that the cost of $i$ with respect to $T$ comes from a route in which they take the bus, i.e., there exist stops $t_{1},t_{2}\in T$ such that $i$ walks from $\ell_{i}$ to $t_{1}$, takes the bus to $t_{2}$, and then walks to $r_{i}$. If $i\notin C_{t_{1}}\cup C_{t_{2}}$, then there exist $s_{1},s_{2}\in S$ such that $s_{1}$ is between $\ell_{i}$ and $t_{1}$, and $s_{2}$ is between $t_{2}$ and $r_{i}$. It is easy to see that $c_{i}(T)=c_{i}(\{t_{1},t_{2}\})\geq c_{i}(\{s_{1},s_{2}\})\geq c_{i}(S)$, contradicting the cost improvement of $i$ according to $T$. Hence, it follows that $i\in C_{t_{1}}\cup C_{t_{2}}$.
We conclude that
[…]
This is a contradiction, and hence, $S$ is in the $2$-core.
To prove that the bound is tight, let $0<\epsilon\leq 1$. We define an instance that is parameterized by an integer $k$. First, observe that $\frac{k+1}{2k}(2-\epsilon)$ converges to $1-\frac{\epsilon}{2}$ as $k$ tends to infinity. Hence, we can choose $k$ large enough so that $\frac{k+1}{2k}(2-\epsilon)\leq 1-\frac{\epsilon}{4}$. Moreover, fix an integer $x>\frac{4}{\epsilon}$. We define an instance with $b=2k$ and $k$ groups of agents $N^{1},\dots,N^{k}$ of size $x$ each, so $n=kx$. We want the agents’ left and right terminals to be separated by a sufficiently long part of the path with no agent terminals. To this end, for each $j\in[k]$, we let $\ell^{j}=j$, $r^{j}=2k+(2j-1)k$, and set
  • $\ell_{i}:=\ell^{j}$ for each $i\in N^{j}$;
  • $r_{i}:=r^{j}$ for x1x-1 agents in $N^{j}$ and $r_{i}=r^{j}+k$ for the remaining agent in $N^{j}$.
@@
For $V=\cup_{i\in N}\{\ell_{i},r_{i}\}$, the algorithm outputs the $2k$ bus stops in $S=[k]\cup\{2k+2jk\colon j\in[k]\}$.
@@
$|M|$ $=n-k=k(x-1)>k\left(x-\frac{\epsilon}{4}x\right)=\left(1-\frac{\epsilon}{4}\right)kx$
$\geq\frac{k+1}{2k}\left(2-\epsilon\right)kx=\left(k+1\right)\left(2-\epsilon\right)\frac{n}{b}\text{.}$
Here, the first inequality holds by our choice of $x$, and the second inequality holds by our choice of $k$. Consider the set of bus stops
$T=\{k\}\cup\{2k+(2j-1)k\colon j\in[k]\},$
with $|T|=k+1$. To show that $M$ is a blocking coalition for the $(2-\epsilon)$-core, it remains to argue that for every agent $i\in M$ it holds that $c_{i}(T)<c_{i}(S)$. To see this, fix a $j\in[k]$ and $i\in M\cap N^{j}$. Then, given solution $T$, $i$ can walk to $k$ and take the bus all the way to $r_{i}$. However, in $S$, $i$ can board the bus at $\ell_{i}$, but then she needs to walk at least $k$ to her destination. Since the overall distance she travels in the latter case is at least as large as in the former case, and involves strictly more walking ($k$ instead of kjk-j), agent $i$ prefers $T$ to $S$. ∎
We note that $\alpha$-BSP can be shown to be a special case of committee selection with monotonic preferences and uniform costs. Hence, the results of Jiang et al. (2020) (Theorem 1) imply that the $16$-core of $\alpha$-BSP is always non-empty. However, Theorem 4.5 offers a much better approximation guarantee.
It remains an open problem how to construct solutions in the core; in fact, we do not even know if the core is always non-empty. This seems to be a challenging question. For instance, by Proposition 2.5, the cost-minimal solution does not even provide JR. Similarly, the solution that selects the most popular terminal points may not provide JR; an example is given in Appendix A.
@@ 4.2 Computation of Outcomes in the Core on Synthetic Data
While it remains open whether the core of $\alpha$-BSP is always non-empty, our simulations indicate that Algorithm 1 can frequently find outcomes in the core. This is despite the fact that, according to Proposition 4.4 and Theorem 4.5, in the worst case this algorithm may fail to find JR outcomes for $\alpha\in(0,1)$ and outcomes in the $\beta$-core for $\beta<2$. Computations were performed using an Apple M2 CPU with 24 GB of RAM. All source code is publicly available at https://github.com/latifian/Bus_Stop_Model.
@@ 4.2.1 Experimental Setup
Figure 5: A heat map showing the frequency of core violations (as a percentage, i.e., 0.05 means 0.05%) of the solutions computed by Algorithm 1 for pairs of $m$ and $b$ in instances of 0-BSP. Each cell is averaged over all values of $n$.
@@
We measure the frequency of instances on which Algorithm 1 computes solutions that provide JR/are in the core. For comparison, we benchmark our algorithm against the naive algorithm, which selects the stops as if the agents were distributed uniformly on the line. More precisely, if $V$ is the set of possible stops, the naive algorithm ignores the actual agents’ types and assumes instead that for each point $p\in\{1,\ldots,100\}$, $n/50$ agents have one of their endpoints at $p$ and report a point in $\arg\min d(p,V)$ as the respective terminal; it then runs Algorithm 1 under this assumption on the agents’ types. In other words, Algorithm 1 can be seen as a weighted version of the naive algorithm, which takes actual user demands into account. In Section C.3, we consider a second benchmark algorithm, which is based on the idea of selecting the most demanded bus stops.
@@ 4.2.2 Experimental Results
The primary goal of our experiments is to measure the performance of Algorithm 1 in terms of achieving outcomes in the core or providing JR beyond the guarantee of Theorem 4.2. Figure 5 shows an aggregated view of our results. We see that, for a fixed $\alpha$, Algorithm 1 computes a solution in the core in more than $99.9\%$ of the cases. Even for the parameter combinations with the highest frequency of fairness violations, the failure rate of Algorithm 1 with respect to computing solutions in the core does not exceed 3%. Figure 5 gives a glimpse at a more nuanced distribution of the failures of computing solutions in the core. We defer a more detailed analysis to Section C.2. We remark that the upper left triangle of the figure cannot contain any core violations because then the budget exceeds the number of stops. One trend that can be observed from Figure 5 and that is confirmed in our analysis for any fixed number of agents is that the frequency of core violations is the highest for a comparatively high number of potential bus stops and a smaller budget.
Another interesting observation is that, for a significant fraction of the instances, if the solution computed by Algorithm 1 is not in the core, it fails JR as well; indeed, this fraction tends to increase as the cost parameter $\alpha$ increases. Moreover, Algorithm 1 performs much better than the naive algorithm. On average it performs up to $230.2$ better for smaller values of $\alpha$ and still $83.7$ times better for $\alpha=0.9$. A detailed comparison of the performance of the two algorithms is provided in Table 1 in Section C.2. Our interpretation of this finding is that, while placing bus stops uniformly is a simple and appealing approach, taking into account the actual user demands results in much fairer solutions.